The formula used to compute a large-sample confidence interval for p is

± (z critical value)

 
(1 − )
n

What is the appropriate z critical value for each of the following confidence levels? (Round your answers to two decimal places.)

(a)    95%

(b)    90%

(c)    99%

(d)    80%

(e)    80%
Answer #1

Using normal table \(P\left(Z>z_{\alpha / 2}\right)\)

a) Here \(\alpha=0.05\) The value of \(z\) is \(1.96\)

b) Here \(\alpha=0.10\) The value of \(z\) is \(1.65\)

c) Here \(\alpha=0.01\) The value of \(\mathrm{z}\) is \(2.58\)

d) Here \(\alpha=0.20\) The value of \(z\) is \(1.28\)